OpenStudy (anonymous):

Find the x- and y-intercepts for the following equation. y = 2x + 4

OpenStudy (haseeb96):

for x intercept put y=0 in the given equation 0=2x+4 2x=-4 x=-4\2 x=-2 and for y intercept put x=0 in the given equation y=2(0)+4 y=4